Name: Astrid Ludwig

Content and Structure – The paper

1. Has a clearly stated topic or question that it will address – 10

2. Does address the topic or question – 5

3. Has a clearly articulated and logical structure:

a. First paragraph lays out the topic or question – 10 b. Later paragraphs have topic sentences and clearly identifiable main ideas – 10 c. Transitions between paragraphs are logical – 10

4. Accurately presents facts and viewpoints of sources – 10

5. Provides sufficient textual evidence to support the argument – 10

6. Is concise, clear, and not too repetitive – 10

7. Has a clearly identifiable conclusion – 10

8. Has a convincing argument that draws together the evidence presented – 10

Total: 95 (A)

Style and English Usage – The paper

1. Is clear, direct, and appropriate in tone – 20

2. Is free of misspellings – 10

3. Is free of grammatical mistakes – 15

4. Avoids incomplete sentences – 10

5. Avoids run-on or overly complicated sentences – 9

6. Uses correct punctuation – 10

7. Is free of jargon, cliches, and misused words – 10

8. Cites quotations and/or references correctly – 10

Total: 99 (A)

Comments: Astrid, this is a good improvement over the first version of your essay. All the technical aspects are fine now. The main issues I still have, though, are ones I mentioned in the comments on the first version too.

First, as a matter of principle, couldn’t one say that a doctor or medical researcher should care about the human being the cells she or he works with came from? I don’t doubt that the way you describe things is probably the way that George Gey and others did feel about HeLa once they had the cells. But is that the way we would or should want to be treated if our tissues are used in experiments? I guess one could argue that doctors must have some distance from the people they treat or else they might not be able to do what is necessary in some cases. But on the other hand, part of the modern version of the Hippocratic Oath does call on doctors to see the human beings they treat as human beings (not diseases or conditions). It seems to me that there’s a case to be made that that ideal should extend to people doing medical research too. Where does the line get drawn?

Second, you really dismiss the possible interest Henrietta’s family might have had in the world’s knowing that HeLa came from a member of their family. Why does that not matter? Yes the cells are not Henrietta once they are removed from her body. But they did come from her body originally.